OpenStudy (anonymous):

The perimeter of a rectangle is 26 feet and the length is 2 foot longer than the width. Find the dimensions an. write a system of a linear equation and solve the resulting system. let x be the length and y be the width. first equation 2x+2y= second equation x+y= what is the length ? what is the width ?

OpenStudy (anonymous):

2x+2y=20? x+y=6 length 8 and width 2 ?

OpenStudy (souvik):

2x+2y=26 x-y=2

Directrix (directrix):

 Also, y = x + 2 So, solve x + y = 13 and y = x + 2 simultaneously. Post what you get and I'll tell you what I got. @bherre12